Paul is twice as old a Justin. In 15 years Paul will be 9 years older than Justin will be. How old is Paul now?
Answer by Paul(988)

Justin's age = x
Paul's age = 2x
In 15 years Paul's age = 15+2x
In 15 years Justin's age = x+15

Expression:
2x+15=x+15+9
x=9

2(9)=18
Hence Paul is 18 years old.
Paul.